\begin{thebibliography}{KMM87}

\bibitem[Ale92]{Alexeev_LogCanSings}
V.~Alexeev, \emph{Log canonical surface singularities: arithmetical approach},
  Flips and abundance for algebraic threefolds, Soci\'et\'e Math\'ematique de
  France, Paris, 1992, Papers from the Second Summer Seminar on Algebraic
  Geometry held at the University of Utah, Salt Lake City, Utah, August 1991,
  Ast\'erisque No. 211 (1992), pp.~47--58.

\bibitem[Ale93]{Alexeev_22Terms}
\bysame, \emph{Two two-dimensional terminations}, Duke Math. J. \textbf{69}
  (1993), no.~3, 527--545.

\bibitem[Bat92]{Batyrev_EffDiv}
V.~V. Batyrev, \emph{The cone of effective divisors of threefolds},
  Proceedings of the International Conference on Algebra, Part 3 (Novosibirsk,
  1989) (Providence, RI), Contemp. Math., vol. 131, Amer. Math. Soc., 1992,
  pp.~337--352.

\bibitem[Ara05]{Ara05}
C.~Araujo, \emph{The cone of effective divisors of log varieties after Batyrev} math.AG/0502174 (2005), Preprint.

\bibitem[Fuj05]{Fujino_Termination}
O.~Fujino, \emph{On termination of 4-fold semi-stable log flips}, Publ. Res.
  Inst. Math. Sci. \textbf{41} (2005), no.~2, 281--294.

\bibitem[Fuj06]{Fuj06}
\bysame, \emph{Special termination and reduction theorem}
Preprint.

\bibitem[HM05]{HaconMcKernan_Flips}
C.~Hacon and J.~McKernan, \emph{On the existence of flips}, arXiv
  math.AG/0507597 (2005), Preprint.

\bibitem[Kaw88]{Kawamata_Crepant}
Y.~Kawamata, \emph{Crepant blowing-up of {$3$}-dimensional canonical
  singularities and its application to degenerations of surfaces}, Ann. of
  Math. (2) \textbf{127} (1988), no.~1, 93--163.

\bibitem[Kaw92]{Kawamata_Term3logflips}
\bysame, \emph{Termination of log flips for algebraic {$3$}-folds},
  Internat. J. Math. \textbf{3} (1992), no.~5, 653--659.

\bibitem[Kaw03]{Kawamata_Term4Flips}
\bysame, \emph{Termination of log flips in dimension 4}, arXiv
  math.AG/0302168v5 (2003), Preprint (withdrawn).

\bibitem[K+92]{Kea92}
J.~Koll{\'a}r et al., \emph{Flips and abundance for algebraic threefolds.}
Soci\'et\'e Math\'ematique de France, Paris 1992. Paper from the Second Summer Seminar on Algebrai Geometry held at the University of Utah, Salt Lake City, Utah, August 1991, Ast\'erisque No. 211 (1992).

\bibitem[KM98]{KollarMori_Book}
J.~Koll{\'a}r and S.~Mori, \emph{Birational geometry of algebraic varieties},
  Cambridge Tracts in Mathematics, vol. 134, Cambridge University Press,
  Cambridge, 1998, With the collaboration of C. H. Clemens and A. Corti,
  Translated from the 1998 Japanese original.

\bibitem[KMM87]{KMM}
Y.~Kawamata, K.~Matsuda, and K.~Matsuki, \emph{Introduction to the minimal
  model problem}, Algebraic geometry, Sendai, 1985, Adv. Stud. Pure Math.,
  vol.~10, North-Holland, Amsterdam, 1987, pp.~283--360, also
  available at {\tt http://faculty.ms.u-tokyo.ac.jp/\~{}kawamata/index.html.}  

\bibitem[Kol89]{Kollar_Flops}
J.~Koll{\'a}r, \emph{Flops}, Nagoya Math. J. \textbf{113} (1989), 15--36.

\bibitem[Mat91]{Matsuki_Term4Flops}
K.~Matsuki, \emph{Termination of flops for {$4$}-folds}, Amer. J. Math.
  \textbf{113} (1991), no.~5, 835--859.

\bibitem[Mum61]{Mumford_SurfaceSingularities}
D.~Mumford, \emph{The topology of normal singularities of an algebraic surface
  and a criterion for simplicity}, Inst. Hautes \'Etudes Sci. Publ. Math.
  (1961), no.~9, 5--22.

\bibitem[Sho85]{Shokurov_Nonvanishing}
V.~V. Shokurov, \emph{A nonvanishing theorem}, Izv. Akad. Nauk SSSR Ser. Mat.
  \textbf{49} (1985), no.~3, 635--651.

\bibitem[Sho03]{Sho03}
\bysame, \emph{Prelimiting flips}. Tr. Mat. Inst. Steklova, 240 (Biratsion. Geom. Linein. Sist. Konechno Porozhdennye Algebry):82-219 (2003).

\bibitem[Sho04]{Shokurov_Letters5}
\bysame, \emph{Letters of a bi-rationalist. {V}. {M}inimal log discrepancies
  and termination of log flips}, Tr. Mat. Inst. Steklova \textbf{246} (2004),
  no.~Algebr. Geom. Metody, Svyazi i Prilozh., 328--351.

\end{thebibliography}

\documentclass[]{amsart}
\usepackage{bm}

\newcommand\defn[1]{{\bf #1}} 
\newcommand{\inv}{^{-1}}

\newcommand{\wZ}{\widetilde Z}
\newcommand{\wB}{\widetilde B}
\newcommand{\wC}{\widetilde C}

\newcommand{\bN}{\mathbb N}
\newcommand{\bZ}{\mathbb Z}
\newcommand{\bQ}{\mathbb Q}
\newcommand{\bR}{\mathbb R}
\newcommand{\bC}{\mathbb C}

\newcommand{\dra}{\dashrightarrow}

\newcommand{\dw}{{d}}
\DeclareMathOperator{\ndiff}{\operatorname{diff}}
\DeclareMathOperator{\Blowup}{\operatorname{Bl}}

\DeclareMathOperator{\codim}{\operatorname{codim}}
\DeclareMathOperator{\Exc}{\operatorname{Exc}}
\DeclareMathOperator{\Supp}{\operatorname{Supp}}
\DeclareMathOperator{\Sing}{\operatorname{Sing}}

\def\Div{\operatorname{Div}}

\newcommand\Q{{\mathbb{Q}}}
\newcommand\R{{\mathbb{R}}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{conditions}[theorem]{Conditions}
\theoremstyle{definition}

\newtheorem{definition}[theorem]{Definition}
\newtheorem{remark}[theorem]{Remark}

\newtheorem{example}[theorem]{Example}

\newtheorem*{acknowledgements}{Acknowledgments}

\author{Valery Alexeev}
\address{Department of Mathematics, University of Georgia, Athens GA
  30602, USA}

\author{Christopher Hacon}
\address{Department of Mathematics, University of Utah, Salt Lake
  City, UT 84112-0090, USA}

\author{Yujiro Kawamata}
\address{Department of Mathematical Sciences, University of Tokyo,
  Komaba, Meguro, Tokyo, 153-8914, Japan}  

\title{Termination of (many) 4-dimensional log flips}
\date{May 3, 2006}

\begin{document}
\begin{abstract}
  We prove that any sequence of 4-dimensional log flips that begins with
  a klt pair $(X,D)$ such that $-(K_X+D)$ is numerically equivalent to
  an effective divisor, terminates.  This implies termination of flips
  that begin with a log Fano pair and termination of flips in a
  relative birational setting.  We also prove termination of directed
  flips with big $K_X+D$.
  As a consequence, we prove existence of minimal models of
  4-dimensional dlt pairs of general type, existence of 5-dimensional
  log flips, and rationality of Kodaira energy in dimension 4.
\end{abstract}
\maketitle

Let $$(X,D) = (X^0,D^0)\dra (X^1,D^1)\dra (X^2,D^2)\dra \dots$$ be a
sequence of flips, relative over another variety $Z$, such that
\begin{enumerate}
\item $X^n$ are normal complex varieties of dimension 4, and $D^n=\sum
  d_jD_j^n$ are $\bR$-divisors,
\item $D^n$ are birational transforms of $D$ on $X^n$,
\item the pair $(X,D)$ is klt (Kawamata log terminal).
\end{enumerate}

The purpose of this note is to prove the termination of such sequences
in several situations. One of them is the case when $-(K_X+D)$ is
numerically equivalent over $Z$ to an effective $\bR$-divisor $E$,
i.e. $-(K_X+D)-E$ is an $\bR$-Cartier divisor which is zero on
curves contracted by $X\to Z$ (Theorem \ref{thm:negative-K}).  Note
that this effectivity condition holds in any of the following
situations:

\begin{enumerate}
\item $-(K_X+D)$ has nonnegative Kodaira dimension,
\item $X\to Z$ is generically finite. Indeed, in this case any
  divisor, including $-(K_X+D)$, is big over $Z$.
\end{enumerate}

Another situation is when $(X,D)$ is of general type over $Z$,
i.e. $K_X+D$ is big over $Z$. In this case we prove that a sequence of
\emph{directed} flips terminates (Theorem~\ref{thm:general-type}).
This allows us to prove that every 4-dimensional dlt pair of general type has a
minimal model (Corollary~\ref{cor:general-type}), and that
5-dimensional log flips exist (Theorem~\ref{5dim-flips}).
The paper is organized as follows:

In Section \ref{sec:preliminaries} we briefly review some
prerequisites.

Section~\ref{sec:Creative counting} is the heart of the paper. There,
we introduce a new \emph{difficulty} of a klt pair, improving on the
methods that were previously used by Shokurov, Kawamata, Matsuki,
Fujino and others \cite{Shokurov_Nonvanishing, KMM, Matsuki_Term4Flops,
  Kawamata_Term3logflips, Fujino_Termination, Kawamata_Term4Flips,
  Shokurov_Letters5}. Some key ideas for our definition come
from the unpublished manuscript \cite{Kawamata_Term4Flips} of 
Kawamata (using homology groups of the boundary divisors to store
information) and from the paper \cite{Shokurov_Letters5} of Shokurov
(using weights).  

A counting argument allows us to prove that for any fixed
$\alpha<1$ no divisors with discrepancy equal to $\alpha$ can appear
over flipping or flipped locus infinitely many times.
Then in Section~\ref{sec:Termination} we derive from this statement
the termination of flips in various useful situations. In the case
when $-(K_X+D)$ is effective, the termination is achieved in one
step. When $K_X+D$ is big, we first prove that sequences of
directed flips terminate. Then we establish the existence of a minimal
model. 

Section~\ref{sec:Applications} lists several consequences that follow
from our results. These include existence of 5-dimensional log flips
and the rationality of Kodaira energy for 4-dimensional klt pairs.

\begin{acknowledgements}
  A large part of this paper was written during the first author's
  2004 visit to RIMS, Kyoto University. He would like to express his
  gratitude for the hospitality and a very stimulating atmosphere. He
  was also partially supported by NSF grant 0401795.
  The second author was partially supported by NSF grant
  0456363 and by an AMS centennial scholarship.  
\end{acknowledgements}

\section{Preliminaries}
\label{sec:preliminaries}

We work over $\bC$ and use standard facts and notations of the Minimal
Model Program, as in \cite{KollarMori_Book}. Let us briefly recall
some of them.

Let $(X,D)$ be a pair consisting of a normal variety $X$ and an
$\bR$-Weil divisor $D=\sum d_jD_j$. It will be a standing assumption
that the divisor $K_X+D$ is $\bR$-Cartier. Then for every divisorial
valuation $v$ of the function field of $X$ one has the
\defn{discrepancy} (not the log discrepancy!)  $a(v, X,D)$ or simply
$a_v$ with respect to $K_X+D$.  The \defn{center of valuation}
$v$ on a variety $W$ will be denoted by $C(v,W)$. If $F=C(v,W)$ is a
divisor then we may write $a(F)$ for $a_v$.  For the components $D_j$
of $D$ one has $a(D_j, X,D) = -d_j$.

A pair $(X,D=\sum d_jD_j)$ is
\begin{enumerate}
\item \defn{terminal} if $d_j<1$ and for all $v\ne v_{D_j}$ one has
$a(v,X,D)>0$; we say that the
  pair is \defn{effective} if $d_j\ge 0$.
\item \defn{klt} (Kawamata log terminal) if $0<d_j<1$ and all $a(v,X,D)>-1$.
\end{enumerate}

A \defn{flip}, relative over $Z$, is a diagram
$
X^- \overset{\psi^-}{\longrightarrow} W
\overset{\psi^+}{\longleftarrow} X^+
$
of birational transformations of varieties over $Z$ such that
\begin{enumerate}
\item $\psi^-$ and $\psi^+$ are proper birational morphisms which are
  isomorphisms outside of closed subsets of codimension $\ge2$,
\item the relative Picard ranks of $X^-$ and $X^+$ over $W$ equal one,
\item $-(K_{X^-} + D)$ is $\psi^-$-ample and $K_{X^+} + D$ is $\psi^+$-ample.
\end{enumerate}
(Whenever two varieties are isomorphic in
codimension 1, we may denote birational transforms of divisors by the
same letters, for convenience.)
If $X^-$ is $\bQ$-factorial then so is $X^+$.  The exceptional set
$\Exc\psi^-$ is called the \defn{flipping locus} and $\Exc\psi^+$ the
\defn{flipped locus}.  

The most basic and well-known fact about flips is this:
\begin{lemma}\label{lem:flip-improves}
  Let $ X^- \overset{\psi^-}{\longrightarrow} W
  \overset{\psi^+}{\longleftarrow} X^+ $ be a flip. Then for any
  divisorial valuation $v$ one has $a(v,X^-,D^-) \le a(v,X^+,D^+)$ and
  the inequality is strict precisely for $v$ with $C(v,X^-) \subset
  \Exc\psi^-$, equivalently with $C(v,X^+)\subset \Exc\psi^+$.
\end{lemma}
In particular if $(X^-,D^-)$ is klt then so is $(X^+,D^+)$.

\begin{lemma}[\cite{KMM}, Lemma 5.1.17]
  \label{lem:types-of-flips}
  $\dim\Exc\psi^- + \dim\Exc\psi^+ \ge \dim X^- -1$. 
In particular, for $\dim X=4$ the possibilities for the pair 
$(\dim\Exc\psi^- , \dim\Exc\psi^+)$ are $(2,1)$, $(2,2)$ and $(1,2)$.
\end{lemma}

A \defn{terminalization} of a klt pair $(X,D)$ is a terminal pair
$(Y,B)$ together with a morphism $f:Y\to X$ such that $K_Y+B =
f^*(K_X+D)$. For any $v$ one has $a(v, X,D) = a(v, Y,B)$. All the
terminalizations used in this paper will be $\bQ$-factorial varieties.

\begin{lemma}\label{lem:terminalization}
  A $\bQ$-factorial terminalization exists for any klt pair $(X,D)$
  with $\dim X=4$.  
\end{lemma}
\begin{proof}
  For the proof, see \cite[Thm.5]{Kawamata_Term3logflips}. The proof
  is valid in any dimension provided the existence of flips (which for
  4-dimensional flips was proved by Shokurov and Hacon-McKernan
  \cite{Sho03,HaconMcKernan_Flips}) and the termination of terminal flips
  (Fujino \cite{Fujino_Termination}).
%
\end{proof}

A klt pair $(X,D)$ has only finitely many divisorial valuations with
$a_v\in (-1,0]$. They are:
\begin{enumerate}
\item the components $D_j$ of $D$, and 
\item the exceptional divisors of a terminalization $(Y,B)\to (X,D)$.  Let
$\bold{e(X,D)}$ denote the \defn{number of these exceptional
  divisors}.
\end{enumerate}
By Lemma~\ref{lem:flip-improves} in a sequence of flips the number
$e(X^n,D)$ 
decreases, and is eventually constant. 
There are more valuations in the next interval $(0,1)$:

\begin{example}\label{ex:echos}
  Let $(Y,B=\sum b_iB_i)$ be a terminal pair, and let $C\subset Y$ be a
  codimension-2 subvariety lying on a unique irreducible component
  $B_i$ of $B$ and not contained in $\Sing B_i$.
  Since $Y$ is terminal, it is nonsingular along the
  generic point of $C$.
  Let $E_1$ be the unique irreducible divisor of the blowup 
  $Y_1=\Blowup_{C}\overset{\pi}{\longrightarrow} Y$
  which dominates $C$, and let $C_1= E_1 \cap \pi_*\inv B_i$, the
  latter denoting the strict preimage of $B_i$. Let $E_2$ be the
  unique irreducible divisor of the blowup
  $Y_2=\Blowup_{C_1}Y_1$, etc. Then a simple computation (which is the
  same as in the case of a surface $Y$ and a point $C$) gives
  \begin{displaymath}
    a(E_1)= 1-b_i, \ a(E_2)= 2(1-b_i), \dotsc , \
    a(E_k)= k(1-b_i)
  \end{displaymath}
  We will call the divisorial valuation corresponding to $E_k$ the
  \defn{$\bold{k}$-th echo} of $Y$ along~$C$. 
\end{example}

\begin{lemma}\label{lem:echos}
  Let $(Y,B)$ be a terminal pair. Then the only divisorial valuations
  with $a(v, Y,B)\in (0,1)$ are:
  \begin{enumerate}
  \item The echos of $Y$ along irreducible subvarieties $C$ of
    codimension 2 that lie on a unique $B_j$, as above. For each such
    $C$, only finitely many $k$-th echos satisfying $k(1-b_j)<1$
    appear.
  \item Finitely many others.
  \end{enumerate}
\end{lemma}
\begin{proof}
  The proof is an easy application of the formula (see, e.g.,
  \cite{Kollar_Flops}) computing the discrepancies of a log smooth
  pair, applied to a log resolution of $(Y,B)$ on which the finitely
  many divisors $F_j$ with $a(Y,B,F_j)\le 0$ are disjoint.
\end{proof}


\begin{lemma}\label{lem:exc-divs-independent}
  Let $\pi:X\to Y$ be a birational morphism of projective varieties 
  with smooth $X$ and normal $Y$, and let
  $E_1,\dotsc,E_m$ be the irreducible components of the exceptional
  locus with $\codim_X E_i =1$. Then
  \begin{enumerate}
  \item $E_i$ are linearly independent in the Neron-Severi group $N^1(X)$
  \item The integer
    $$
    {\rho(Y)} = \dim N^1(X)\otimes\bR -
    \#(\text{of exceptional divisors } E_i)
    $$
    is well defined, i.e. does not depend on a particular resolution
    of $Y$, and is positive. 
  \end{enumerate}
\end{lemma}
\begin{proof}
  (1) Suppose that there is a nontrivial linear relation, and let $e$
  be the maximal dimension of $\pi(E_i)$ that appear. By cutting with
  $e$ general hyperplanes on $Y$ and $\dim X - e+2$ hyperplanes on $X$
  we can reduce to the case when $X$ and $Y$ are surfaces. By
  \cite{Mumford_SurfaceSingularities} the intersection form on the
  exceptional curves $E_i$ is negative definite, so no nontrivial
  linear relation between $E_i$ is possible.

  (2) Let $\phi:X'\to X$ be another resolution of singularities, with
  exceptional divisors $F_i$.  Then we have 
  homomorphisms $\phi_*$ and $\phi^*$ between the Neron-Severi groups.
  One has $\phi_*\phi^* D_1=D_1$ and $\phi^*\phi_*D_2 = D_2 + \sum n_j
  F_j$ for some $n_j\in \bZ$.  Therefore, $\dim N^1(X') -
  \dim N^1(X) = \#(F_j)$, and so $\rho(Y)$ is well defined. It is
  positive because $E_i$ are linearly independent and do not span
  $N^1(X,\bR)$. 
\end{proof}


\begin{remark}
  Our final remark is about dlt (divisorially log terminal) flips. A
  well-known statement called Special Termination (see e.g.
  \cite[2.1]{Fuj06}) reduces termination of dlt flips to the klt
  flips.  So all our termination results can be strengthened from klt
  to dlt.
\end{remark}


\section{Creative counting}
\label{sec:Creative counting}

Ever since Shokurov gave the original argument in the 3-dimensional
terminal case \cite{Shokurov_Nonvanishing}, the main idea for proving
termination of flips has remained the same. One introduces an appropriate
\emph{difficulty} $d(X,D)$ which counts valuations with certain
discrepancies, and then shows that $d(X^-,D^-) > d(X^+,D^+)$ for a
flip. Then one can conclude the termination if the difficulty is, say,
a nonnegative integer, or under somewhat weaker assumptions.

The main approach in this paper will be the same, with several twists.
As Example~\ref{ex:echos} and
Lemma~\ref{lem:echos} show, there are infinitely many valuations in $(0,1)$,
but except for the echos, there are only finitely many of them. The
new ideas are these: 
\begin{enumerate}
\item Count discrepancies in $(0,1)$ with a normalizing summand, so that the
  echos contribute zero to the sum, thus making it finite.
\item Count them with weights (cf. \cite{Shokurov_Letters5}).
\item Use the Picard groups of the boundary divisors $B_i$ on a
  terminalization $(Y,B)$ of $(X,D)$ appropriately to ``store'' some
  information (cf.  \cite{Kawamata_Term4Flips}).
\end{enumerate}

We first define the \defn{difficulty of a terminal pair}. That will be
a real number $\dw(X,D) $ which will depend on a 
\defn{weight function} $w:(0,+\infty)\to \bR_{\ge0}$ on the
set of discrepancies. The weight function is assumed to satisfy the following:
\begin{conditions}\label{conditions}
  \begin{enumerate}
  \item $w(a)\ge 0$ for all $a$, and $w(a)=0$ for $a\ge1$. 
  \item $w$ is decreasing (when we say \defn{increasing} or
    \defn{decreasing}, we always mean \defn{not necessarily strictly}).
  \item \label{item4b}
    Whenever $0< \sum m_i b_i < 1$ for some $m_i\in \bZ_{\ge0}$,
    one has
    $$w(1-\sum m_i b_i) \ge \sum m_i w(1-b_i).$$
  \item For $i=1,\dotsc, m$, let  $\{b_i^n,\ n\in\bN\}$ be $m$
    decreasing sequences of real numbers in $(0,1)$. Then there exists
    $n_0$ such that the following holds:

    Whenever $k\ge 2$, $m_i\ge0$ and $n\ge n_0$ are integers such that
    \begin{displaymath}
      1 - \sum m_i b_i^n \in (0,1), \quad\text{resp.}\quad
      k(1-b_1^n) - \sum_{i>1} m_i b_i^n  \in (0,1) 
    \end{displaymath}
    then one has
    \begin{displaymath}
      \ndiff_1(n) \ge \ndiff_1(n+1)  \quad\text{resp.}\quad
      \ndiff_2(n) \ge \ndiff_2(n+1),
    \end{displaymath}
    where
    \begin{align*}
&    \ndiff_1(n) = w(1 - \sum m_i b_i^n) - \sum m_i w(1-b_i^n) 
    \quad \text{and} \\
&    \ndiff_2(n) = 
    w\big( k(1-b_1^n) -  \sum_{i>1} m_i b_i^n \big) - w \big( k(1-b_1^n) \big)
    \end{align*}
  \end{enumerate}
\end{conditions}

For example, we could take for $w$ a piecewise linear function with
decreasing absolute values of slopes.
The main two functions we will use in this papers will be
$w_{\alpha}^-$ and $w_{\alpha}^+$, defined as follows:
\begin{definition}\label{def:functions}
  Let $\alpha\in (0,1)$. Then
  \begin{enumerate}
  \item $w_{\alpha}^-(x)=1-x$ for $x\le \alpha$ and
    $w_{\alpha}^-(x)=0$ for $x> \alpha$, and
  \item $w_{\alpha}^+(x)=1-x$ for $x<\alpha$ and $w_{\alpha}^+(x)=0$
    for $x\ge \alpha$.
  \end{enumerate}
  The reader is invited to do an elementary exercise now and check
  that both $w_{\alpha}^-$ and $w_{\alpha}^+$ satisfy the above
  conditions. The integer $n_0$ is chosen so that for $n\ge n_0$ for
  every $i$ the sequence $\{b_i^n\}$ settles into one of the open
  intervals between the points
  \begin{displaymath}
    0, \ 1-\alpha, \ 1-\frac{\alpha}{2},\ \dotsc\
    ,1-\frac{\alpha}{k},\ \dotsc
  \end{displaymath}
  or into one of these points.
\end{definition}

We will define another, \defn{summed weight} function $W:(-\infty,1)\to
\bR_{\ge0}$ by the formula
$$
W(b) = \sum_{k=1}^{\infty} w\big( k(1-b) \big). $$ It easily follows
that there are only finitely many nonzero terms in this sum, that
$W$ is a nonnegative and increasing function, and that $W(b)=0$ for
$b\le 0$.  The meaning of $w$ and
$W$ and the necessity of the above conditions will be clear from the
proofs of Lemmas \ref{lem:nonnegative}, \ref{lem:monotonicity1},
\ref{lem:monotonicity2}.

Let $\nu:\coprod \wB_i \to \cup B_i$ be the normalization of the
divisor $\Supp B$. For any irreducible subvariety $C\subset \Supp B$
the preimage $\wC = \nu\inv(C)$ splits into a union of irreducible
components $\wC_{ij}\subset\wB_i$.

\begin{definition}\label{defn:diff-terminal}
  Let $(Y,B)$ be a terminal pair, \emph{not necessarily effective,}
  i.e. possibly with some $b_i<0$. Then we define
  \begin{eqnarray*}
   \dw(Y,B) = 
   && \sum_{a(B_j)\le 0} W(b_i) \rho( \wB_i ) + 
   \sum_{v; \ \codim C(v,Y)\ne 2 } 
   w( a_v ) + \\
   && +\sum_{ {\rm irr.} C \subset Y; \codim C =2}
   \left[
     \sum_{v; \ C(v,Y)=C } 
     w( a_v ) - \sum_{\wC_{ij}} W(b_i) 
   \right]
 \end{eqnarray*}
\end{definition}

\begin{remark}
  For $w(x)=\max(1-x,0)$ this definition is a version of Shokurov's
  ``stringy'' invariant $\rho^2$ in \cite{Shokurov_Letters5} except
  that we take into account higher ``echos'' (which is necessary).
\end{remark}

\begin{lemma}\label{lem:well-defined}
  $\dw(Y,B)$ is well defined, i.e. only finitely many summands are
  nonzero.
\end{lemma}
\begin{proof}
  Indeed, if $C$ is a codimension-2 subvariety of $Y$ lying on a
  unique component $B_i$ and not in $\Sing B_i$, the contribution
  coming from the echo divisors of $C$ is zero. And by
  Lemma~\ref{lem:echos} there are only finitely many other nonzero terms.
\end{proof}

\begin{lemma}\label{lem:nonnegative}
  Assume that $(Y,B)$ is an effective terminal pair. Then 
  $\dw(Y,B) \ge 0$.
\end{lemma}
\begin{proof}
  Clearly, $\sum W(b_i) \rho( \wB_i )\ge 0$ and $w(a_v)\ge0$, so we
  only need to prove that the contributions from codimension-2
  subvarieties are nonnegative. This is a computation that can be done
  assuming that $B_i$'s are curves on a nonsingular surface $Y$
  passing through a nonsingular point $C$ with multiplicities $m_i$.
  Let $E_1$ be the exceptional divisor of the blowup of $Y$ at $C$. We
  have $a(E_1)= 1-\sum m_ib_i$. Let us order the divisors so that
  $b_1\ge b_2 \ge\dotsc $.

  \emph{Case 1: all $b_i<1/2$.} Then $W(b_i)=w(1-b_i)$. 
  Let $m_i'$ be the number of
  irreducible components $\wC_{ij}$ lying over $C$; then 
  $m_i' \le m_i$. Therefore, we obtain
  \begin{align*}
    \sum_{v;\ C(v,Y)=C } 
    w( a_v ) - \sum_{\wC_{ij}} W(b_i) \ge
    & \ w(1-\sum m_ib_i) - \sum m'_i w(1-b_i) \\
    \ge & \  w(1-\sum m_ib_i) - \sum m_i w(1-b_i) 
  \end{align*}
  which is nonnegative by Condition~\ref{conditions}(3).

  \emph{Case 2: $b_1\ge 1/2$.} Then, since $(Y,B)$ is terminal, we
  must have $m_1=1$ and $b_i<1/2$ for $i\ge2$. Then in the negative
  direction we have additional contributions coming from $W(b_1)$, and
  they are $w( k(1-b_1))$ for all $k\ge 2$.

  On the other hand, let $E_2$ be the exceptional divisor of the
  blowup at the intersection point of $E_1$ and the strict
  preimage of $B_1$; $E_3$ be the exceptional divisor of the blowup
  at the intersection point of $E_2$ and the strict preimage of $B_1$, etc. 
  Continuing this way by induction, for each $k\ge 2$ we obtain a
  divisor $E_k$ with
  $$
  a(E_k) =k(1-b_1) -\sum_{i>1} m_i^k b_i
  \quad \text{for some } m_i^k\ge 0.
  $$
  Since the function $w$ is decreasing by Condition
  \ref{conditions}(2), we have
  \begin{displaymath}
    w\big( a(E_k) \big) 
     \ge  w\big( k(1-b_1)\big).
  \end{displaymath}
  So the positive additional terms outweigh the negative ones.
\end{proof}

\begin{lemma}\label{lem:pullback}
  Let $f:Y'\to Y$ be a proper birational morphism and $(Y',B')$,
  $(Y,B)$ be two terminal pairs such that $K_{Y'}+B' = f^*(K_{Y}+B)$.
  Then $\dw(Y',B') = \dw(Y,B)$.
\end{lemma}
\begin{proof}
  Let $\wB'_i \to \wB_i$ be the induced birational morphisms between
  the normalizations of the divisors with $a(v)\le 0$. The only
  difference between $\dw(Y',B')$ and $\dw(Y,B)$ is in the
  codimension-2 subvarieties $\wC'_{ij}\subset \wB'_i$ whose image on
  $\wB_i$ has codimension $\ge3$. They appear in the first and in the
  last terms, and they cancel out by Lemma~\ref{lem:exc-divs-independent}. 
\end{proof}

\begin{definition}
  The \defn{difficulty of a klt pair} $(X,D)$ is defined to be 
  $\dw(X,D) = \dw(Y,B)$, where $f:(Y,B)\to (X,D)$ is any terminal pair
  such that $K_Y+B = f^*(K_X+D)$. 
\end{definition}

\begin{lemma}\label{lem:nonnegative2}
  If $(X,D)$ is a klt pair then $\dw(X,D)\ge 0$. 
\end{lemma}
\begin{proof}
  Indeed, we can use for $(Y,B)$ a terminalization of $(X,D)$, and
  then the statement follows by Lemma~\ref{lem:nonnegative}.
\end{proof}

\begin{lemma}[Monotonicity 1] \label{lem:monotonicity1}
  Let $(Y^1,B^1)$ and $(Y^2,B^2)$ be two birational terminal pairs
  that have the same ``$a(v)<0$'' parts, i.e. 
  for every component $B_i^1$ with $b_i^1>0$ there exists a
  birational to it component $B_i^2$ with $b_i^2=b_i^1$, and vice
  versa. Assume that one has
  \begin{displaymath}
    (\varphi^1)^*(K_{Y^1}+B^1) \ge (\varphi^2)^*(K_{Y^2}+B^2) 
  \end{displaymath}
  on a common log resolution $\varphi^n: U \to Y^n$. Then
  $\dw(Y^1,B^1) \ge \dw(Y^2,B^2)$.
\end{lemma}
\begin{proof}
  By Lemma \ref{lem:pullback} we can replace both pairs by their
  pullbacks on $U$. Then the only difference is between $w(a_v)$, and
  it is nonnegative because the weight function is decreasing. 
\end{proof}

\begin{lemma}[Monotonicity 2]\label{lem:monotonicity2}
  For $i=1,\dotsc, m$ let $\{b_i^n,\ n\in\bN\}$ be $m$ decreasing
  sequences of real numbers in $(0,1)$, and let $n_0$ be chosen as in
  Condition~\ref{conditions}(4).  Let $(Y,B^n=\sum b_i^n B_i)$ be the
  effective terminal pairs with the same variety $Y$.  Then for $n\ge
  n_0$ one has $\dw(Y,B^n)\ge \dw(Y,B^{n+1})$.
\end{lemma}
\begin{proof}
  We clearly have the required inequality for the first term $\sum
  W(b_i)\rho(\wB_i)$. Also, for all valuations with $\codim_Y
  C(v,Y)>2$ one has $a(v,Y,B^n) \le a(v,Y, B^{n+1})$ and so for the
  corresponding terms in Definition \ref{defn:diff-terminal} one gets
  $w(a(v,Y,B^n)) \ge w(a(v,Y, B^{n+1}))$.

  It remains to consider valuations with $\codim_Y C(v,Y)=2$. As in
  the proof of Lemma~\ref{lem:nonnegative}, we have
  \begin{enumerate}
  \item the terms 
    \begin{align*}
      &    w(1 - \sum m_i b_i^n) - \sum m_i w(1-b_i^n) 
      \quad \text{and} \\
      &    
      w\big( k(1-b_1^n) -  \sum_{i>1} m_i^k b_i^n \big) 
      - w\big( k (1-b_1^n) \big),
    \end{align*}
  \item   and then some other positive terms of the form 
    $w(c - \sum m''_i b_i^n)$ for some $m''_i\ge 0$. (We used the fact
    that $m'_i \le m_i$.)
  \end{enumerate}
  For the terms in (1) we have the required inequality by
  Condition~\ref{conditions}(4). For the terms in (2) the inequality
  follows from the fact that $w$ is a decreasing function.
\end{proof}


\begin{theorem}\label{thm:klt-flips-decrease}
  Let $(X,D) = (X^0,D)\dra (X^1,D)\dra (X^2,D)\dra \dots$ be a
  sequence of flips with $(X,D)$ klt.  Then there exists $n_0$ such
  that for $n\ge n_0$ the sequence $\dw(X^n,D)$ is decreasing.
\end{theorem}
\begin{proof}
  Let $ X^n \overset{\psi^-}{\longrightarrow} W
  \overset{\psi^+}{\longleftarrow} X^{n+1} $ be the $n$-th flip.  For
  $n$ large enough, we have $e(X^n,D) = e(X^{n+1},D)$, i.e. on the
  terminalizations $(Y^n,B^n)$, $(Y^{n+1},B^{n+1})$ the irreducible
  components of $B^{n+1}$ are birational transforms of the irreducible
  components of $B^n$. Then we have sequences $\{b_i^n\}$ of
  coefficients of $B^n$, and we choose $n_0$ large enough, as in
  Lemma~\ref{lem:monotonicity2}. 

  Let $\varphi^-:U_n\to Y^n$, $\varphi^+:U_n\to Y^{n+1}$ be a variety
  dominating both terminalizations. Denoting the birational transform
  $(\psi^+_*)\inv \psi^-_* B^n$ on $Y^{n+1}$ again by $B^n$, we have
  $B^n \ge B^{n+1}$ on $Y^{n+1}$ by Lemma~\ref{lem:flip-improves}.
  Further, we claim that
  \begin{displaymath}
    (\varphi^-)^* (K_{Y^n} + B^n) \ge 
    (\varphi^+)^* (K_{Y^{n+1}} + B^n) 
  \end{displaymath}
  Indeed, the first divisor is numerically negative over $W$, and
  therefore also over $Y^{n+1}$, and the second divisor is zero over
  $Y^{n+1}$.  Hence, the difference is negative over $Y^{n+1}$, and it
  consists of exceptional divisors of $\varphi^+$. This implies that
  the difference is effective.

  Now, we have $\dw(Y^n, B^n) \ge \dw(Y^{n+1}, B^n)$ by
  Lemma~\ref{lem:monotonicity1} and $\dw(Y^{n+1}, B^n) \ge \dw(Y^{n+1},
  B^{n+1})$ by Lemma~\ref{lem:monotonicity2}.
\end{proof}

To summarize what we have done this far: For every weight function $w$
satisfying Conditions~\ref{conditions} and for any klt pair $(X,D)$ we
have defined difficulty $\dw(X,D)\ge 0$ and proved that in a sequence
of flips it decreases at each step for $n\ge n_0$. Now, if we could
only establish that it decreases by at least a fixed $\epsilon>0$, that would
certainly imply termination. 
So far we have not used the fact that for some valuations $v$, the
weight $w(a_v)$ may drop abruptly. We only used the fact that after a flip one
has $a_v(X^-,D^-) \le a_v(X^+,D^+)$ and so $w(a_v(X^-,D^-)) \ge
w(a_v(X^+,D^+))$.

\begin{theorem}\label{thm:fixed-A}
  Fix a number $\alpha\in (-1,1)$. Then in a sequence of flips 
  $$(X,D) = (X^0,D^0)\dra (X^1,D^1)\dra (X^2,D^2)\dra \dots$$ 
  there cannot be infinitely many $k$ for which there exist valuations
  $v$ with $a_v=\alpha$ whose center is in the flipping or the flipped
  locus.
\end{theorem}
\begin{proof}
  If $\alpha\in (-1,0]$, the situation is easy, since there are only
  finitely many discrepancies $a_v\in(-1,0]$ and after a flip $a_v$
  strictly increases. 
  
  So assume $\alpha\in (0,1)$.  For the flipping locus we use the
  function $w_{\alpha}^-$ of \ref{def:functions}. Then after a flip
  the corresponding discrepancy changes from $a_v(X^n,D^n)=\alpha$ to
  $a_v(X^{n+1},D^{n+1})>\alpha$.  Hence, the weight
  $w_{\alpha}^-(a_v)$ and the difficulty drop down by $(1-\alpha)$.
  This can happen only finitely many times. For the flipped locus we use
  the function $w_{\alpha}^+$ instead.
\end{proof}

\begin{lemma}\label{lem:no-sing-codim2}
  In a sequence of klt flips it cannot happen infinitely many times
  that the flipping or flipped locus has a component which has
  codimension $2$ in $X^n$ and is contained in the singular
  locus of $X^n$.
\end{lemma}
\begin{proof}
  Let $E_1, \dots, E_m$, $m=e(X^n,D)$, be the finitely many valuations
  of $X^n$ with $a(v,X^n,D)\le 0$ for $n\ge n_0$.

  Assume that $X^n$ is generically singular along a codimension-2
  component. By cutting with two generic hyperplanes we obtain a log
  terminal singular surface pair $(Z,D|_Z=\sum d_j D_j|_Z)$, where
  $D_j|_Z$ need no longer be irreducible. Let $\wZ$ be the minimal
  resolution of singularities, with exceptional curves $F_k$. Then for
  each $F_i$, $a(F_i,Z,D|_Z) \le 0$, and so $a(F_i,Z,D|_Z) = a(E_i,
  X^n,D)$ for some $E_i$. Since there are only finitely many $E_i$'s,
  there will be a subsequence such that for one of the divisors, say
  $E_1$, the discrepancy $a(E_1,X,D)$ appears as the minimal
  discrepancy of $(Z,D|_Z)$, and increases on every step. But the set
  of minimal discrepancies of log terminal pairs satisfies the
  ascending chain condition by \cite[Thm.3.2,3.8]{Alexeev_22Terms}.
  Contradiction.

  \emph{Second proof}. The positive numbers $-F_i^2$ are bounded from
  above by $2/(1+ \min(a_v(X,D)))$. If we could bound the number of
  $F_i$'s, there would be only finitely many possibilities for the
  weighted graphs of minimal resolutions of $Z$'s, and so finitely
  many possible indexes, and finitely many possible discrepancies
  $a(F_i)$. The statement then would easily follow.  

  Each divisor $E_i$ can lead to many $F_i$'s, so we cannot limit the
  number of curves on the minimal resolution of $Z$. However, it is
  still true that the number of \emph{distinct} discrepancies
  $a(F_i,Z,D|_Z)$ is $\le m$. Now, the dual graph of the minimal
  resolution of a singularity on $Z$ is a tree with at most one fork
  and three legs
  (by the classification of log terminal surface singularities, see
  e.g. \cite{Kawamata_Crepant,Alexeev_LogCanSings}).
  A basic computation shows that on the interior of each leg
  \begin{enumerate}
  \item the function $a(F_i)$ is concave up, and
  \item if $a(F_{i-1})=a(F_i)=a(F_{i+1})$ then $F_i^2=-2$ and $F_i\cdot D|_Z=0$.
  \end{enumerate}
  Hence, the only way one can have repeating discrepancies in the
  interior of a leg is when you have chains of $(-2)$-curves. These
  chains can be shortened without changing the set of
  discrepancies. Hence, we can assume that in the interior of each leg
  you have no more than 2 repeating $a(F_i)$'s. We are reduced to the
  finitely many graphs with at most $3m+7$ vertices, and so we are
  done by the above argument.
\end{proof}

\begin{theorem}\label{thm:no-codim2}
  In a sequence of klt flips it cannot happen infinitely many times
  that the flipping or flipped locus has a component which has
  codimension $2$ in $X^n$ and which is contained in $D^n$.
\end{theorem}
\begin{proof}
  Let $C$ be such a codimension-2 component.  By the previous lemma we
  can assume that $Y$ is nonsingular generically along $C$. The blowup
  along $C$ produces a divisor with $a_v = 1-\sum m_ib_i <1$. We have
  a finite set of such discrepancies in $(-1,1)$, and we apply
  Theorem~\ref{thm:fixed-A} to each of them to complete the proof.
\end{proof}

\section{Termination and minimal models}
\label{sec:Termination}

\begin{lemma}\label{lem:positive-bdry}
  Let $$(X,D) = (X^0,D)\dra (X^1,D)\dra (X^2,D)\dra \dots$$ be a
  sequence of 4-dimensional klt flips such that $D=E + D'$ with
  effective $E$ and $D'$, such that all flips are $E$-positive (i.e.
  $E$ is positive on a curve in the flipping locus).  Then the
  sequence of flips terminates.
\end{lemma}
\begin{proof}
  Write $E=\sum e_jD_j$ as the sum of irreducible components, with
  $e_j>0$. Then for every flip there exists a component $D_j$ such
  that $D_j$ is positive on the flipping locus, and so is negative on
  the flipped locus. So, $D_j$ must contain the flipped locus. 

  By Theorem~\ref{thm:no-codim2} after some $n\ge n_0$ the flipped
  locus does not contain any codimension-2 components. By
  Lemma~\ref{lem:types-of-flips} this implies that all the flips for
  $n\ge n_0$ are the $(2,1)$-flips. But for each of these flips the
  dimension of the homology group $H_4(X^n,\bR)$ goes down, and this
  cannot be repeated infinitely many times.
\end{proof}

\begin{theorem}\label{thm:negative-K}
  Let $$(X,D) = (X^0,D)\dra (X^1,D)\dra (X^2,D)\dra \dots$$ be a
  sequence of klt flips, relative over another variety $Z$.
  Assume that $-(K_X+D)$ is numerically equivalent over $Z$ to an effective
  $\bR$-divisor $E$ 
  (i.e. the difference is an $\bR$-Cartier divisor which is
  zero on curves contracted by $X\to Z$).
   Then the sequence of flips is finite.
\end{theorem}
\begin{proof}
  Note that $E\ne0$. 
  Replace $D$ by $D+ \epsilon E$. Then for $0<\epsilon \ll1$ the pair
  $(X,D+\epsilon E)$ is still klt and since $K+D+\epsilon E \equiv
  (1-\epsilon)(K+D)$, the sequence of flips for $K+D$ is a sequence of
  flips for $K+D+\epsilon E$. Since $E$ is positive on each flipping
  locus, we are done by the previous Lemma~\ref{lem:positive-bdry}.
\end{proof}

Our next application is to the case when $K_X+D$ or $D$ are big.  Let
us begin by recalling the following consequence of the Cone Theorem
(which follows immediately from \cite{KMM}):

\begin{lemma}\label{L1} 
  Let $(X,D)$ be a klt pair and $f:X\to Z$ a projective morphism to a
  normal variety.  Let $A$ be an effective $\R$-Cartier divisor such that
  $K_X+D$ is not relatively nef, but $K_X+D+A$ is relatively nef. Then
  there is a $K_X+D$ extremal ray $R$ over $Z$ and a real number
  $0<\lambda \leq 1$ such that $K_X+D+\lambda A$ is relatively nef,
  but trivial on $R$.

  If $D$ and $A$ are $\bQ$-Cartier divisors then $\lambda\in \bQ$.
\end{lemma}

Given $(X,D)$ as above and an ample line bundle $A$ on $X$, one may
run the \defn{MMP for $\bold{(X,D)}$ directed by $\bold{A}$.} 
Recall here that 4-dimensional log flips exist by Shokurov and 
Hacon-McKernan \cite{Sho03,HaconMcKernan_Flips}.
We proceed as follows: Since
$A$ is ample, then $K_X+D+tA$ is nef for some $t\gg 0$. Let
\begin{displaymath}
  t_0=\inf \{ t\geq 0 | K_X+D+tA\ {\rm is\  relatively\ nef }\}.
\end{displaymath}
If $t_0=0$, or $R$ induces a Mori-Fano fibration, we stop,
otherwise we replace $(X,D)$ by the pair $(X^1, D)$ given by flipping
or contracting $R$.  Now $K_{X^1}+D+t_0A$ is relatively nef and by
\eqref{L1} we can repeat the above trick.  At each step, we obtain a
pair $(X^n, D)$ and a real number $t_n$ such that $K_{X^n}+D+t_{n}A$
is relatively nef and $t_n$ is the infimum of all $t\geq 0$ such that
$K_{X^n}+D+tA$ is relatively nef.  Note that $X^{n+1}$ is obtained
from $X^n$ by performing a $(K_{X^n}+D)$-flip or divisorial contraction
with extremal ray $R_n$.

It is clear that each flip or divisorial contraction is $(K+D)$-negative
and $A$-positive and we have the following possibilities:
\begin{enumerate}
\item $K_{X^n}+D$ is relatively nef;
\item $R^n$ induces a relative Mori-Fano fibration for $K_{X^n}+D+t_nA$ and
  hence for $K_{X^n}+D$;
\item For all $n\geq n_0$, $R^n$ induces a $(K_{X^n}+D)$-flip.
\end{enumerate}
  
\begin{theorem}\label{thm:general-type}
  Let $(X,D)$ be a 4-dimensional klt pair over $Z$.
  Assume that for some $c_j\in \bR$ the $\bR$-divisor $c_0 K_X + \sum
  c_jD_j$ is numerically equivalent to a big divisor.
  Then there exists a finite sequence of flips and divisorial
  contractions $X^n\dasharrow X^{n+1}$, $0\leq k\leq n$
  (here $X^0=X$) such that $K_{X^n}+D$ is nef 
  or there exists a $K_{X^n}+D$ Mori-Fano fibration.
\end{theorem}
\begin{proof}
  Denote $D'=\sum c_jD_j$. Write the big divisor as $A+E$ with an
  effective divisor $E$ and an ample divisor $A$ whose support does
  not have components in common with $D+E$. Then
    \begin{eqnarray*}
&    K+D \equiv 
    (1-\epsilon c_0)K + D-\epsilon D' + \epsilon A + \epsilon E
    =
    (1-\epsilon c_0)(K + D''), \\
&    \text{where}\quad
    D'' = (1-\epsilon c_0)\inv (D-\epsilon D' + \epsilon A + \epsilon E)
    \end{eqnarray*}

For $0<\epsilon\ll 1$ the pair $(X,D'')$ is klt, the divisor $D''$ is
effective and contains a positive multiple of $A$. If we run the MMP
directed by $A$ then each extremal ray will be $A$-positive. So by
Lemma~\ref{lem:positive-bdry} the flips terminate. So we either arrive
to a minimal model of $K+D''$, and hence of $K+D$ which is
proportional to it, or we end up with a Mori-Fano fibration.
\end{proof}

\begin{corollary}\label{cor:dltgt}
Let $(X,D)$  be a $4$-dimensional $\Q$-factorial dlt pair, such that
$D\in \Div _\Q(X)$ is big. Let $A$ be any ample divisor on $X$.
Then the MMP for $(X,D)$ directed by $A$ terminates.
\end{corollary}
\begin{proof} It suffices to show that there is no infinite sequence of 
$K_X+D$ flips (directed by $A$) $(X^n,D)\dasharrow (X^{n+1},D)$.
By \cite[2.1]{Fuj06}, we know that after finitely many flips, the flipping 
and flipped loci are disjoint from the locus of log canonical singularities 
$\lfloor D\rfloor $. We now conclude as in the proof of Theorem
\ref{thm:general-type}. 
\end{proof}

\begin{corollary}\label{cor:general-type}
  Let $(X,D)$ be a dlt pair of general type over $Z$. Then there exists
  a minimal model of $(X,D)$ over $Z$.
\end{corollary}
\begin{proof}
  Indeed, the Mori-Fano fibration in this case is not possible. 
\end{proof}

\section{Further applications}
\label{sec:Applications}

We begin by recalling a result concerning the big cone of a
klt pair.
Let $(X,D)$ be a projective $\Q$-factorial klt pair, $D\in \Div_\Q (X)$ and $A$ an ample divisor on $X$.
Then one has:
\begin{definition} If $K_X+D$ is not pseudo-effective, then 
the \defn{effective log-threshold} of $A$ is defined by
$$\sigma (X,D,A)={\operatorname {sup}}\{t\in \Q \ |\ A+t(K_X+D)\ {\rm is\ effective}\}.$$ The \defn{Kodaira Energy} is defined by $\kappa\epsilon (X,D,A)=-1/\sigma (X,D,A)$.\end{definition}
Following the approach of Batyrev \cite{Batyrev_EffDiv}, we deduce the
following result concerning the rationality of the Kodaira Energy (see
\cite{Ara05} for more details).
\begin{theorem}\label{thm:KE} Let $(X,D)$ be a projective $4$-dimensional
$\Q$-factorial klt pair and $A$ an ample divisor on $X$.
If $K_X+D$ is not pseudo-effective, then $\sigma (X, D , A)$ is rational.
\end{theorem}
\begin{proof} 
  We use the MMP for $(X,D)$ directed by $A$.  Since $K_X+D$ is not
  pseudo-effective, it ends with a Mori fiber space.  Let $R^n$ be the
  final extremal ray on the model $(X^n,D)$.  Let $\lambda$ be the
  nef threshold which is known to be rational.  Then
  $K_{X^n}+D+\lambda A^n$ is nef, hence pseudo-effective, and
  $(K_{X^n}+D+\lambda A^n) \cdot R^n=0$.  Since $-(K_{X^n}+D)$ is
  relatively ample for the Mori fiber space, we conclude that 
  $\sigma(X,D,A)=-1/\lambda\in \bQ$ as required.
%
%

\end{proof}
We now turn our attention to the problem of the existence of flips in
dimension~$5$. We have the following:

\begin{theorem}\label{5dim-flips}
  Let $(X,D)$ be a $5$-dimensional klt $\Q$-factorial pair, $D\in \Div
  _\Q (X)$ (or $\Div_\R(X)$), $f:(X,D)\to Z$ a flipping contraction.
  Then the flip $f^+:X^+\to Z$ of $f$ exists.
\end{theorem}
In \cite{HaconMcKernan_Flips}, the existence of flips in dimension $n$
is claimed assuming the termination of flips in dimension $n-1$.
Since we have not shown the termination of flips for $4$-folds in full
generality, we must analyze the given argument in more detail. 
\begin{proof} By \cite{HaconMcKernan_Flips}, PL-flips exist in dimension $5$,
provided that flips (relative over a variety $Z$) 
terminate for any klt pair $(T, B)$ where $T\to Z$ is birational and
$\dim T=4$. This is the case by Theorem \ref{thm:negative-K}.
We now follow Shokurov's Reduction Theorem as outlined in \S 3 of 
\cite{Fuj06}.
We do not reproduce the entire argument, but simply indicate the changes 
required to 3.7 of \cite{Fuj06}.

In Step 2 of 3.7, we have
$\nu:Y\to X$ an appropriate log resolution of $X$ and 
a morphism $h:Y\to X\to Z$. 
$H'$ is an appropriately chosen divisor on $Z$ and $H:=f^*H'$.
We may assume that $D$ contains a general $f$-ample $\Q$-divisor $D'$. We let
$A\sim _{\Q} e_0\nu ^*D'+\sum e_iE_i$ be a general ample divisor over $Z$ 
where $0<e_0\ll 1$, $0<-e_i\ll 1$ and $E_i$ are
$\nu$-exceptional divisors. 
Note that $K_Y+(D+H)_Y+\sum e_iE_i\sim_\Q K_Y +(D^0+H)_Y+A$ where $D^0=D-e_0D'\geq 0$ and 
$(D+H)_Y$ denotes the birational transform of $D+H$, 
that is the strict transform of $D+H$
plus the sum of all the exceptional divisors taken with multiplicity $1$.
Notice that we may replace $D$ by $D-e_0D'+\nu _*A$ so that
$K_Y +(D^0+H)_Y+A-\nu ^*(K_X+D)$ is given by the strict transform of $H'$ plus an exceptional divisor.

In Step 3 we run the MMP with respect to $K_Y+(D^0+H)_Y+A$ over $Z$ directed by $A$. All flips are PL-flips and their
termination follows by the usual special termination argument (as in \S 2 of \cite{Fuj06})
noting that to run this argument, it suffices to establish termination of directed flips for dlt pairs with big boundary divisor
in dimension $\leq 4$. This in turn follows from 
Corollary \ref{cor:dltgt}.
We obtain a $\Q$-factorial dlt pair $\bar{h}:(\bar{Y},(D^0+H+A)_{\bar {Y}})\to Z$ such that $K_{\bar{Y}}+(D^0+H+A)_{\bar{Y}}$ is $\bar{h}$-nef.

In Step 5, we subtract $\hat {H}$ (where $\hat{H}$ denotes
the strict transform of $H$ on $\bar{Y}$).
The main point is that for this step to work, we need the special termination
of dlt
flips directed by $\hat {H}$. By the arguments of Special Termination (\S 2 of
\cite{Fuj06}), we need the termination of dlt
flips directed by the restriction $\hat {H}$ for log pairs of dimension
at most $4$ whose boundary is a big $\Q$-divisor. 
This follows as in Corollary \ref{cor:dltgt}.
Therefore we obtain $\tilde{h}:(\tilde{Y},D_{\tilde {Y}}^0+\tilde{A})\to Z$ such that $\tilde {Y}$ is
$\Q$-factorial, $(\tilde{Y},D_{\tilde{Y}}^0+\tilde{A})$ is dlt and $K_{\tilde{Y}}
+D_{\tilde{Y}}^0+\tilde{A}$ is $\tilde{h}$-nef.
By the negativity lemma, one can check that $\tilde{h}$ is small
so that $(\tilde{Y},D_{\tilde{Y}}^0+\tilde{A})$ is in fact klt.
\footnote{To see this, let $W$ be a common log resolution of $\tilde{Y}$ and $X$ with morphisms $\phi :W\to \tilde{Y}$ and $\psi:W\to X$. Then 
$-B:=\phi ^*(K_{\tilde{Y}}+D^0_{\tilde {Y}}+\tilde{A})-\psi ^*(K_X+D)$ is exceptional and $\psi$-nef, therefore $B$ is effective and $\phi ^*(K_{\tilde{Y}}+D^0_{\tilde {Y}}+\tilde{A})+B=\psi ^*(K_X+D)$. If $E\subset \tilde{Y}$ is a divisor which is $Z$-exceptional, then this implies that $a(E,X,D)\leq -1$ which is absurd. It also follows that the canonical models over $Z$ for $(X,D)$ and $(\tilde{Y},D^0_{\tilde {Y}}+\tilde{A})$ coincide.}
One sees that the canonical model of 
$(\tilde{Y},D_{\tilde{Y}}^0+\tilde{A})$ over $Z$ coincides with the
canonical model of $(X,D)$ over $Z$ and this is 
the required flip.
\end{proof}

\bibliographystyle{amsalpha}

\def\cprime{$'$}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}